### v1

# On the breakdown point of transport-based quantiles

Abstract
Recent work has used optimal transport ideas to generalize the notion of (center-outward) quantiles to dimension $d\geq 2$. We study the robustness properties of these transport-based quantiles by deriving their breakdown point, roughly, the smallest amount of contamination required to make these quantiles take arbitrarily aberrant values. We prove that the transport median defined in Chernozhukov et al. (2017) and Hallin et al. (2021) has breakdown point of $1/2$. Moreover, a point in the transport depth contour of order $\tau\in[0,1/2]$ has breakdown point of $\tau$. This shows that the multivariate transport depth shares the same breakdown properties as its univariate counterpart. Our proof relies on a general argument connecting the breakdown point of transport maps evaluated at a point to the Tukey depth of that point in the reference measure.
Keywords: Center-outward quantiles, contamination model, multivariate medians, optimal transport, robustness, Tukey depth.
MSC2020 subject classifications. Primary: 62G35, 62G30.
1 Introduction
Recently, [10, 23] proposed a new notion of multivariate quantiles based on mass transportation. They naturally lead to multivariate definitions of ranks and signs that are distribution free and can be used to construct statistically efficient nonparametric tests [23, 13]. The idea behind transport quantiles is to preserve two fundamental properties of the quantile function $Q_{X}$ of a univariate random variable $X\sim P$: it pushes the uniform (Lebesgue) measure on $[0,1]$ forward to $P$; and it is monotone, i.e., it is the derivative of a convex function. The transport map provides a natural multivariate analogue construction. Given a random vector $X\sim P$ and a reference measure $\mu$, the transport-based quantile function is the unique (a.s. defined) gradient of a convex function $\mathbf{Q}_{\pm}$ pushing $\mu$ forward to $P$. The existence and uniqueness of $\mathbf{Q}_{\pm}$ is guaranteed by McCann’s theorem [36].
The univariate empirical median is one the most fundamental examples of a robust statistic. It estimates consistently the population median and it is highly robust in the presence of outliers. In particular, the minimum fraction of bad sample values required for the median to take arbitrarily large values is $1/2$. In other words, its breakdown point is $1/2$. In this paper we show that the breakdown point of the multivariate transport-based median is also $1/2$. More generally, we show that the transport-based depth contour of order $\tau$ is defined as the image of the Tukey depth contour of $\mu$ via the transport-based quantile map. We show that the transport-based depth contour of order $\tau\in[0,1/2]$ has breakdown point $\tau$. This general result is analogous to the breakdown point of the empirical univariate quantile of order $q=k/n\in(0,1)$ which is exactly $\min(k,n+1-k)/n$. Note that the breakdown points that we find are dimension independent and they are attained by computationally efficient estimators, unlike many classical high breakdown methods in the robust statistics literature [35, 12, 18, 46, 49].
The proof of the breakdown point of the transport-based quantile contours relies on an unexpected relation between the optimal transport problem and the Tukey depth. We show that the breakdown point of transport maps evaluated at a point of the support of the reference measure is equal to the Tukey depth of that point in the reference measure. Therefore in order to maximize the breakdown point of transport quantiles on has to choose reference measures $\mu$ that attains the maximal Tukey depth of $1/2$. This is the case of natural choices of $\mu$; c.f. 1.3 and 3.8.
1.1 Breakdown point of the optimal transport problem
It turns out that studying the aforementioned breakdown points of the transport-based quantiles, is equivalent to answering the following question, which itself is of independent interest:
Given a sample $X_{1},\dots,X_{n}$, what is the minimum fraction of bad sample values required to make the optimal transport map from $u_{1},\dots,u_{n}$ to $X_{1},\dots,X_{n}$ evaluated at a particular $u_{i}$ arbitrarily bad?
Before delving into details, let us recall the definition of the optimal transport problem for discrete measures, We refer to [47] for an in-depth exposition on the topic. Let $X^{(n)}=\{X_{1},\dots,X_{n}\}$ be a sample in $\mathbb{R}^{d}$ (usually assumed to follow a distribution $P$) and $u^{(n)}=\{u_{1},\dots,u_{n}\}\subset\mathbb{R}^{d}$ be a fixed deterministic sequence. We call $u^{(n)}$ and $X^{(n)}$ the reference and target samples, respectively. Call $\Gamma(u^{(n)},X^{(n)})$ the set of all of bijective mappings $T:u^{(n)}\to X^{(n)}$. An empirical optimal transport map $\hat{T}_{n}=\hat{T}_{X^{(n)}}$ is any solution of the (quadratic) optimal transport problem
$\min_{T\in\Gamma(u^{(n)},X^{(n)})}\frac{1}{n}\sum_{i=1}^{n}\|T(u_{i})-u_{i}\|^{2}.$ (1)
Answering our main question is equivalent to determining the finite sample breakdown point of the estimator $\hat{T}(u_{i})$, i.e.,
${\rm BP}(\hat{T}(u_{i}))=\frac{1}{n}\min\left\{\ell\in\{1,\dots,n\}:\ \sup_{Z^{(n)}\in\mathcal{Q}_{\ell,n}}\|\hat{T}(u_{i})-\hat{T}_{Z^{(n)}}(u_{i})\|=\infty\right\},$
where $\mathcal{Q}_{\ell,n}$ is the set of all sets $Z^{(n)}=\{Z_{1},\dots Z_{n}\}\subset\mathbb{R}^{d}$, sharing exactly $n-\ell$ elements with $X^{(n)}$. The breakdown point is one of the simplest and more popular measures of robustness in statistics. The finite sample definition given above introduced by [19] in the context of univariate location estimators and quickly extended to various multivariate settings in the robust statistics literature e.g. [43, 44, 12, 48]. The notion of breakdown point was originally introduced using an asymptotic formulation in a [25, 26]. We will also introduce a definition of breakdown of transport maps for population measures akin to the asymptotic formulation of the breakdown point in Section 2.
Both the median and quantiles are the standard univariate measures of depth, indicating how extreme a point is with respect to a dataset. However, the definitions of median and quantiles, based on the canonical order of the real line, do not naturally extend to multivariate data. In higher dimensions, depth theory offers an alternative. The literature provides various notions of depth, such as Tukey depth, spatial depth, integrated depth, or lens depth (see [39] and references therein). The most commonly used is the Tukey depth . Recall from [45] that the Tukey (or halfspace) depth of a point $x\in\mathbb{R}^{d}$ w.r.t. the set $X^{(n)}$ is defined as
$\mathrm{TD}(x;X^{(n)})=\min_{v\in\mathcal{S}^{d-1}}\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\left(\langle v,X_{i}-x\rangle\geq 0\right)}\right\},$
where $\mathcal{S}^{d-1}=\{v:\|v\|=1\}$ denotes the unit sphere and $\mathbf{1}$ the indicator function, i.e., $\mathbf{1}_{A}(u)=1$ if $u\in A$ and $\mathbf{1}_{A}(u)=0$ otherwise. It takes values in the interval $[0,1/2]$, where the observation with maximum depth is called the Tukey median and is considered the most central, while observations with depth close to 00 are regarded as more extreme. The breakdown point of the Tukey median is $1/3$ for halfspace-symmetric distributions and $1/(d+1)$ for general configurations of points.
In this work we answer the motivating question by finding an unexpected relation between the optimal transport problem and the Tukey depth. The following main result assume the reference sequence of points is in general position, i.e., no more than $d$ points lie in any $(d-1)$-dimensional affine subspace.
Theorem 1.1.
Let $\hat{T}$ be a solution to (1) and assume $u^{(n)}$ is in general position. Then,
${\rm BP}(\hat{T}(u_{i}))\in\left[\mathrm{TD}(u_{i};u^{(n)})-\frac{d-1}{n},\mathrm{TD}(u_{i};u^{(n)})\right].$
The upper bounds stated in Theorem 1.1 are sharp as discussed in 3.2. In Section 3, we provide a slightly more general version of Theorem 3.1 as well as an analogous statement for the transport map of continuous measures.
1.2 Breakdown points of transport-based medians and contours.
The empirical transport-based quantile function $\mathbf{Q}_{\pm}^{(n)}$ of a sample $X^{(n)}$ of $P$ is any mapping solving (1) while $\mathbf{F}_{\pm}^{(n)}=(\mathbf{Q}_{\pm}^{(n)})^{-1}:X^{(n)}\to u^{(n)}$ is the empirical transport-based distribution function. Note that these definitions depend on the choice of $u^{(n)}$. In practice, one could take a random sample from a population reference measure $\mu$. Deterministic choices include taking a “spherical regular grid” [23] or a Halton sequence in a hypercube [13].
The transport-based depth $D_{\pm}(X_{i})$ of an observation $X_{i}$ was defined in [10] as the Tukey depth of $\mathbf{F}_{\pm}^{(n)}(X_{i})$ with respect to $u^{(n)}$, i.e.,
${\rm D}_{\pm}^{(n)}(X_{i})={\rm TD}(\mathbf{F}_{\pm}^{(n)}(X_{i});u^{(n)}).$
So far, the depth function is defined on $X^{(n)}$. In order to extend its definition to the whole $\mathbb{R}^{d}$, [23] proposed to extend $\mathbf{F}_{\pm}^{(n)}$ by considering a continuous cyclically monotone interpolation, i.e., by choosing a convex and differentiable function $\psi_{n}=\psi_{n}(\cdot;X^{(n)}):\mathbb{R}^{d}\to\mathbb{R}$ such that $\nabla\psi_{n}(X_{i})=\mathbf{F}_{\pm}^{(n)}(X_{i})$ for all $i=1,\dots,n$. Equipped with this extension one can define the empirical transport-based depth function of a point $x\in\mathbb{R}^{d}$ as
${\rm D}_{\pm}^{(n)}(x)={\rm D}_{\pm}^{(n)}(x;X^{(n)}):={\rm TD}(\nabla\psi_{n}(x);u^{(n)})$
as well as the empirical empirical transport-based depth contours
${\cal C}_{\pm}^{(n)}(\tau)=\mathcal{C}_{\pm}^{(n)}(\tau;X^{(n)}):=\{y\in\mathbb{R}^{d}:\ {\rm D}_{\pm}^{(n)}(y)=\tau\},\quad\tau\in[0,1/2].$
An empirical transport-based median $m_{\pm}^{(n)}$ of $X^{(n)}$ is a deepest point in the sense
$m_{\pm}^{(n)}\in\operatorname*{arg\,max}_{x\in\mathbb{R}^{d}}{\rm D}_{\pm}^{(n)}(x)$
Our main result Theorem 1.1 provides asymptotically sharp estimates of the finite sample breakdown point of ${\cal C}_{\pm}^{(n)}$, which is defined as
${\rm BP}({\cal C}_{\pm}^{(n)}(\tau))\\ :=\frac{1}{n}\min\left\{k\in\{1,\dots,n\}:\ \sup_{Z^{(n)}\in\mathcal{Q}_{k,n}}d_{H}(\mathcal{C}_{\pm}^{(n)}(\tau;X^{(n)}),\mathcal{C}_{\pm}^{(n)}(\tau;Z^{(n)}))=\infty\right\}.$
Here $d_{H}$ denotes the Hausdorff distance, which is defined as:
$d_{H}(A,B)=\max\left\{\sup_{a\in A}\inf_{b\in B}\|a-b\|,\sup_{b\in B}\inf_{a\in A}\|b-a\|\right\},\quad A,B\subset\mathbb{R}^{d}.$
Note that, as depth contours are generally compact sets, the Hausdorff distance is a natural metric for measuring distances between depth contours. See for example [18, 6] and [15] in the context of Tukey and transport-based depth contours respectively.
Corollary 1.2.
Let $X_{1},\dots,X_{n}$ be a sequence of vectors in $\mathbb{R}^{d}$ with all of the elements being different and let $u^{(n)}$ be in general position. Set $\tau^{*}:=\max_{u\in\mathbb{R}^{d}}{\rm TD}(u;u^{(n)}).$ Then the following holds irrespective of the dataset $X_{1},\dots,X_{n}$,
  1. 1.
    (Breakdown point of the median)
    ${\rm BP}(m_{\pm}^{(n)})\in\left[\tau^{*}-(d-1)/n,\tau^{*}\right].$
  2. 2.
    (Breakdown point of the depth contours) For every $\tau\in[0,\tau^{*})$ with ${\cal C}_{\pm}^{(n)}(\tau)\neq\emptyset$,
    ${\rm BP}({\cal C}_{\pm}^{(n)}(\tau))\in[\tau-(d-1)/n,\tau].$
In Section 3, we provide a slightly more general version of 1.2 that covers the case where the reference sequence is not assumed to be in general position. We also provide a continuous population counterpart that provides an asymptotic breakdown point in the sense of Hampel [26].
Remark 1.3.
The choice of $u^{(n)}$ affects the value of $\tau^{*}$ which in turn affects the breakdown points of the transport-based median and depth contours. If $u^{(n)}$ is such that $\tau^{*}=1/2$, the resulting median preserves the high breakdown point of its univariate counterpart. However, $u^{(n)}$ should not be chosen very “organized” as the general position assumption could be violated. If $u^{(n)}$ is a sample of a distribution $\mu$ absolutely continuous w.r.t. the Lebesgue measure $\mathcal{L}_{d}$ and $n>d$, then the reference measure is in general position with probability $1$. Furthermore, if additionally, $\mu$ has maximum Tukey depth $1/2$, then $u^{(n)}$ has maximal Tukey depth of $1/2+o(1)$ for $d$ fixed as $n\to\infty$. Note that all the choices that have been suggested as reference measures for transport-based quantiles are half-space symmetric and hence they attain the maximal Tukey depth of $1/2$, e.g. the uniform on the hypercube $[-1,1]^{d}$, isotropic Gaussians and the spherical uniform distribution [10, 21, 23, 28].
Remark 1.4.
Part 2 of 1.2 can be viewed as a multivariate analogue of the well-known breakdown point of $\min\{\tau,1-\tau\}$ for empirical $\tau$-quantiles. Note however that in the univariate case, the $\tau$-depth contour is not a standard $\tau$-quantile, but a “center-outward quantile” as described in [23]. Indeed, the left to right orientation of standard quantiles is replaced by a center-outward orientation that is more amenable to be generalized in to dimension $d\geq 2$. As such, the $1/4$-depth contour does not correspond to the $1/4$-quantile but to end-points of the interquantile range, i.e., the $1/4$- and $3/4$-quantiles. Therefore the probability contained inside a $1/4$-contour is $1/2$, and more generally the probability mass contained inside the region defined by a $\tau$-contour is $2\cdot(1-\tau)$.
1.3 Related work
There is a rich literature focusing on deriving the breakdown point of robust multivariate estimators of location and designing high breakdown point methods. It has been well established that while univariate M-estimators defined by a bounded score function have a high breakdown point, many natural multivariate generalizations of them can have a dimension dependent breakdown [35, 19, 29, 46, 34, 30]. Because the definition of the univariate median relies on the canonical ordering of the real line, it can be generalized in multiple natural ways to multivariate settings starting with $d=2$. As a result, numerous definitions of multivariate medians have been studied in the literature. Prominent examples include the geometric median [31, 8, 37, 38], the Tukey median [45, 18, 9], projection depth-based medians [1, 49, 41, 17], among others. The coordinatewise median and the geometric median [31, 8, 37, 38] can be computed efficiently and are known to have high breakdown point of $1/2$ [32]. However they do not correspond to natural notions of depth as discussed in [50, 10]. Perhaps the most popular depth-based median is the Tukey median and is known to have a breakdown point of $1/3$ [18]. Some other depth-based medians attain the maximal breakdown point of $1/2$ [46, 49]. However, we note computing the Tukey median and high breakdown point projection-based estimators is NP hard [3] which significantly reduces their applicability for $d$ moderately large. We note that not only do transport-based medians have a high breakdown point, but they can also be computed in polynomial time with a computational cost of $O(n^{3})$, irrespective of the dimension (cf.  [40]), where $n$ is the sample size. Furthermore, transport-based quantiles can also naturally be used to define multivariate ranks. Such rank statistics are distribution free and have been shown to be useful for constructing statistically efficient nonparametric tests [23, 13].
We note that the influence function is another common way to characterize the robustness properties of an estimator in the robust statistics literature [27, 24]. This approach is commonly referred to as infinitesimal robustness since it aims to quantify the impact of small amounts of contamination in the statistical functional of interest by finding a Gâteaux derivative at a fixed perturbation $\delta_{x}$. In the context of transport maps, to the best of our knowledge, the form of the influence function $\frac{d}{dt}\big{|}_{t=0}T_{\mu\to(1-t)P+t\delta_{x}}$ has not been established. The closest results we found in the literature involves finding the derivative of $\frac{d}{dt}\big{|}_{t=0}T_{\mu\to(1-t)P+tQ}$ when $Q$ is regular enough (see [33] for periodic measures and [22] for the general setting).
1.4 Organization of the paper
The rest of the paper is organized as follows: In Section 2, we introduce the key definitions related to the population transport problem that we will need in this work. In particular, we introduce the general population optimal transport problem, the canonical extension of transport maps, transport-based depth contours and an appropriate notion of breakdown point. In Section 3, we state the main results of the paper in the form of the breakdown point of the empirical and population transport maps evaluated at any point in the support of the reference measure. These results connect the breakdown point of the transport maps to the Tukey depth of the point in the reference measure, showing that transport-based depth contours of order $\tau\in(0,1/2)$ have a breakdown point of $\tau$. Finally, Section 4 contains the proofs of all the results presented in the paper.
2 Preliminaries
In this section we introduce the population optimal transport problem and some important notions that we will require to state our main results.
2.1 Notation
The triple $(\Omega,\mathcal{A},\mathbb{P})$ is the probability space. Let $\mathcal{P}(\mathbb{R}^{d})$ be the set of Borel probability measures over $\mathbb{R}^{d}$ and $\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ be the set of probability measures absolutely continuous w.r.t. the $d$-dimensional Lebesgue measure $\mathcal{L}_{d}$. The topological closure and interior of a set $A\subset\mathbb{R}^{d}$ is denoted as ${\rm cl}(A)$ and ${\rm int}(A)$, respectively, while ${\rm coh}(A)$ denotes the convex hull of $A$. The support of a Borel probability measure $P$, denoted as ${\rm supp}(P)$, is the smallest closed set (in the sense of contention) with $\mathbb{P}$-probability $1$. Let $2^{\mathbb{R}^{d}}=\{A:\ A\subset\mathbb{R}^{d}\}$ and denote by ${\rm dom}(G)$ the domain of a set-valued function $G:\mathbb{R}^{d}\to 2^{\mathbb{R}^{d}}$ i.e., ${\rm dom}(G)$ is the set of all $x$ such that $G(x)\neq\emptyset$. Finally, let $\mathbb{B}=\{x\in\mathbb{R}^{d}:\leavevmode\nobreak\ \|x\|\leq 1\}$.
2.2 Extension of monotone transport maps
[…]
where $(T_{\#}\mu)(\cdot)=\mu(T^{-1}(\cdot))$ denotes the push-forward measure. We call $\nabla\varphi$ optimal transport map even if the second order moments of $\mu$ and $P$ are not finite. The canonical extension of the transport map $\nabla\varphi$ is the subdifferential
$\mathbb{R}^{d}\ni u\mapsto\partial\varphi(u)=\left\{x\in\mathbb{R}^{n}:\ \varphi(z)\geq\varphi(u)+\langle x,z-u\rangle\quad\forall z\in\mathbb{R}^{n}\right\}.$
It holds that $\{\nabla\varphi(u)\}=\partial\varphi(u)$ for $\mu$-a.e. $u\in\mathbb{R}^{d}$, which shows that $\partial\varphi$ is indeed an extension of $\nabla\varphi$. However, $\partial\varphi$ is no longer taking values in $\mathbb{R}^{d}$, but in $2^{\mathbb{R}^{d}}$. That is, $\partial\varphi$ is a set-valued map. A celebrated theorem by Rockafellar (see [42]) states that subdifferentials of convex functions are maximal monotone, i.e.,
$\langle x_{1}-x_{2},u_{1}-u_{2}\rangle\geq 0,$ (2)
for every $u_{i}\in\mathbb{R}^{d}$ and $x_{i}\in\partial\varphi(u_{i})$, for $i=1,2$, and its graph is not strictly contained in the graph of any other set-valued mapping satisfying (2).
Remark 2.1.
We underline that even if $\nabla\varphi$ is the unique gradient of a convex function pushing $\mu$ forward to $P$, the convex function $\varphi$ might not be unique (cf. [36]) so neither its extension $\partial\varphi$. In Section 3 we show that the breakdown point of the canonical extension of the transport maps does not depend on the choice of the extension.
2.3 Transport-based depth contours
We call $\mu\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ the reference measure. The transport-based quantile function is the unique ($\mu$-a.e. defined) gradient of a convex function $\mathbf{Q}_{\pm}$ pushing $\mu$ forward to $P$. The transport-based distribution function $\mathbf{F}_{\pm}$ of $P$ is the $P$-a.e. defined inverse $\mathbf{F}_{\pm}=(\mathbf{Q}_{\pm})^{-1}$ of the transport-based quantile map. We recall that our goal is to find the breakdown point of $\mathbf{Q}_{\pm}(x)$ for $x\in\mathbb{R}$. Let ${\rm dom}(\mathbf{Q}_{\pm})$ denote the domain of $\mathbf{Q}_{\pm}$ and note that in general $\mathbf{Q}_{\pm}$ is well-defined $\mu$-a.e., so that $\mathbf{Q}_{\pm}(x)$ might be undefined for some $x\in\mathbb{R}^{d}$. Hence, even to guarantee that ${\rm dom}(\mathbf{Q}_{\pm})$ contains the support of $\mu$ one needs to impose regularity conditions over $\mu$ and $P$ in order to guarantee that $\mathbf{Q}_{\pm}$ and $\mathbf{F}_{\pm}$ are well-defined everywhere. A standard condition is to assume that $\mu$ and $P$ have densities supported on a convex set locally bounded away from zero and infinite [7, 20]. However, one of the most natural choices of $\mu$, the so-called spherical uniform distribution $\mu_{d}$ has density ${C_{d}}{\|x\|^{1-d}}$ for some dimension-dependent constant $C_{d}>0$ [23]. Since $\mu_{d}$ has a singularity at zero for $d>1$, it violates the aforementioned standard regularity condition. [^1] One can therefore not rely on this notion of regularity for $\mu_{d}$ in order to get well-defined quantile and distribution functions, irrespective of the regularity of $P$. However, far away from zero, one can still show that the quantile map is well-defined for $P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ and has a density supported on a convex set locally bounded away from zero and infinite (cf. [21, 15, 14]). In particular, the median $\mathbf{Q}_{\pm}(0)$, might be undefined for some arbitrarily regular probability measure $P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$, for $d>2$. (Note that for $d=2$ the Monge–Ampère equation have a slightly nicer behavior, see [21].) In [16] this issue is solved by extending $\mathbf{Q}_{\pm}$ in the “canonical way” introduced for transport maps. Specifically, a set-valued quantile function is
$\mathbb{R}^{d}\ni u\mapsto\mathbb{Q}_{\pm}(u):=\partial\phi(u)=\left\{x\in\mathbb{R}^{n}:\ \phi(z)\geq\phi(u)+\langle x,z-u\rangle\quad\forall z\in\mathbb{R}^{n}\right\},$
where $\phi:\mathbb{R}^{d}\to(-\infty,+\infty]$ is any l.s.c.  convex function such that $\mu(\nabla\phi=\mathbf{Q}_{\pm})=1$. This extension allows us to define the transport-based depth of a point $x\in\mathbb{R}^{d}$ w.r.t. a probability measure $P$ as the set
${\rm D}_{\pm}(x)={\rm D}_{\pm}(x;P):={\rm TD}(\mathbb{Q}_{\pm}^{-1}(x);\mu),$
where
$\mathrm{TD}(u;\mu):=\min_{v\in\mathcal{S}^{d-1}}\mu\left(\{z:\langle v,z-x\rangle\geq 0\}\right)$
is the celebrated Tukey depth and ±$\mathbb{Q}_{\pm}^{-1}(u)=\{u\in\mathbb{R}^{d}:x\in\mathbb{Q}_{\pm}^{-1}(u)\}$ is the set-valued inverse of $\mathbb{Q}_{\pm}$. We assume throughout the rest of the paper that the reference measures used to define the quantiles are half-space symmetric. This is a standard assumption in the literature; c.f. 1.3. The transport-based depth contours are defined as
${\cal C}_{\pm}(\tau)=\mathcal{C}_{\pm}(\tau;P):=\{x\in\mathbb{R}^{d}:\ {\rm D}_{\pm}(x)=\tau\},\quad\tau\in[0,1/2]$
and the transport-based regions as
±$\mathcal{R}_{\pm}(\tau)=\mathbf{Q}_{\pm}(\mathcal{R}^{\mathrm{TD}}_{\mu}(\tau)),\quad{\rm for}\ \tau\in(0,1/2).$
transport-based median as
$m_{\pm}(P)=\operatorname*{arg\,max}_{x\in\mathbb{R}^{d}}{\rm D}_{\pm}(x).$
2.4 Breakdown points of population contours
We define the breakdown point of the depth contour ${\cal C}_{\pm}(\tau;P)$ at $P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ as
${\rm BP}({\cal C}_{\pm}(\tau;P)):=\inf\left\{\varepsilon\in(0,1):\ \sup_{Q\in\mathcal{Q}_{\varepsilon}(P)}d_{H}(\mathcal{C}_{\pm}(\tau;P),\mathcal{C}_{\pm}(\tau;Q))=\infty\right\},$
where ${\cal Q}_{\varepsilon}(P)=\{P+\varepsilon(Q-P):\ Q\in\mathcal{P}(\mathbb{R}^{d})\}\cap\mathcal{P}^{a.c.}(\mathbb{R}^{d})$. Note that for $X_{1},\dots,X_{n}\overset{iid}{\sim}P$ our results (1.2 and 3.7) show that ${\rm BP}({\cal C}_{\pm}^{(n)}(\tau))\to\cal{\rm BP}(C_{\pm}(\tau;P))$ almost surely. One can therefore think of ${\rm BP}({\cal C}_{\pm}(\tau;P))$ as the asymptotic breakdown point of the $\tau$-depth contour at the distribution $P$. This is somehow similar to the original definition of breakdown of Hampel [26].
Clearly our definition breakdown point of ${\cal C}_{\pm}(\tau;P)$ only makes sense if the contour sets are compact. This is automatically guaranteed by 2.2, the compactness of the Tukey depth and the definition of $\cal C_{\pm}(\tau).$
Lemma 2.2.
Set $\mu,P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$. Let $\partial\varphi$ be any maximal monotone extension of the optimal transport map $\nabla\varphi$ from $\mu$ to $P$. Then it holds that
$\{u\in\mathbb{R}^{d}:{\rm TD}(u;\mu)>0\}\subset{\rm int}\left({\rm dom}(\partial\varphi)\right).$
As a consequence, for any compact set $K\subset\{u\in\mathbb{R}^{d}:{\rm TD}(u;\mu)>0\}$ it holds that $\partial\varphi(K)=\bigcup_{u\in K}\partial\varphi(u)$ is compact and nonempty.
Proof.
It is well-known that
$\{u\in\mathbb{R}^{d}:{\rm TD}(u;\mu)>0\}\subset{\rm int}({\rm coh}(\operatorname{supp}(\mu))).$ (3)
Next, as $\nabla\varphi$ pushes $\mu$ forward to $P$, it holds that $\mu({\rm cl}({{\rm dom}(\partial\varphi)}))=1$. Then, by definition of ${\rm supp}(\mu)$, it holds that ${\rm supp}(\mu)\subset{\rm cl}({{\rm dom}(\partial\varphi)})$. Since ${\rm cl}({{\rm dom}(\partial\varphi)})$ is convex (see [2, Corollary 21.14]), it holds that
${\rm coh}(\operatorname{supp}(\mu))\subset{\rm cl}({{\rm dom}(\partial\varphi)}),$ (4)
where for a set $A\subset\mathbb{R}^{d}$, ${\rm coh}(A)$ denotes its convex hull. As ${\rm int}({{\rm dom}(\partial\varphi)})\subset{{\rm dom}(\partial\varphi)}\subset{\rm cl}({{\rm dom}(\partial\varphi)})$ and both ${\rm int}({{\rm dom}(\partial\varphi)})$ and ${\rm cl}({{\rm dom}(\partial\varphi)})$ are convex, it holds that
${\rm int}({{\rm dom}(\partial\varphi)})={\rm int}({\rm cl}({{\rm dom}(\partial\varphi)})).$
The result follows as a consequence of the previous display, (3) and (4). ∎
Our proof of the breakdown of ${\cal C}_{\pm}(\tau;P)$ relies on breaking the extended transport map $\partial\varphi$. This is the key workhorse of our argument. Since $\partial\varphi(u)\neq\emptyset$ might not be unique (cf. 2.1), we use a selection functional to define its breakdown point as follows.
${\rm BP}(\partial\varphi(u),P)=\min\left\{\varepsilon\in(0,1):\sup_{\varphi_{\mu\to\nu}\in\Gamma_{\varepsilon}(P)}\inf_{v\in\partial\varphi_{\mu\to\nu}(u)}\|v\|=\infty\right\},$
where $\Gamma_{\varepsilon}(P)$ is the set of l.s.c. convex functions $\varphi_{\mu\to\nu}$ such that $\nabla\varphi_{\mu\to\nu}$ pushes $\mu$ forward to $\nu\in{\cal Q}_{\varepsilon}(P)$. This selection functional defines the breakdown point as the smallest $\varepsilon$ such that the smallest element of $\partial\varphi_{\mu\to\nu}(u)$ explodes, where $\nu\in\cal Q_{\varepsilon}(P)$. It turns out that the stronger requirement that all the elements $\partial\varphi_{\mu\to\nu}(u)$ explode leads to the same breakdown point. See 3.5.
3 Main results
3.1 Breakdown points of the optimal transport map
This section contains a generalization of Theorem 1.1, valid for reference sequences in non-general position, and its population version. We define the lower Tukey depth of $u$ with respect to the dataset $u^{(n)}$.
$\mathrm{TD}^{-}(u;u^{(n)}):=\frac{n+1}{n}-\max_{v\in\mathcal{S}^{d-1}}\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\left(\langle v,u_{i}-u\rangle\geq 0\right)}\right\}.$
The following result characterizes the breakdown point of the empirical transport problem (1).
Theorem 3.1.
Let $X^{(n)}=\{X_{1},\dots,X_{n}\}$, $u^{(n)}=\{u_{1},\dots,u_{n}\}$ and $\hat{T}$ be as above. Then the breakdown point of $\hat{T}(u_{i})$ lies in between the lower Tukey depth and the Tukey depth of $u_{i}$ with respect to the dataset $u^{(n)}$, i.e.,
${\rm BP}(\hat{T}(u_{i}))\in[\mathrm{TD}^{-}(u_{i};u^{(n)}),\mathrm{TD}(u_{i};u^{(n)})].$
Remark 3.2.
As illustrated in Figure 1 the lower bound in Theorem 1.1 is attained for some reference measures. The upper bound is also attained when we consider the reference and target datasets both equal to $\{1,2,3,4,5\}\subset\mathbb{R}$. The Tukey depth of $3$ is $3/5$ and the breakdown point of the optimal transport map evaluated at $3$ is $3/5$. As a consequence, the bounds of Theorem 1.1 are sharp.
Remark 3.3.
We also observe that $\mathrm{TD}^{-}(u_{i};u^{(n)})$ and $\mathrm{TD}(u_{i};u^{(n)})$ get closer as $n\to\infty$ if $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{u_{i}}$ converges weakly to a probability measure that is absolutely continuous w.r.t. the Lebesgue measure. We underline that the breakdown point of the optimal transport map could be arbitrary small. Indeed, if the reference dataset $u^{(n)}$ is supported on a lower dimensional affine subspace and $X^{(n)}$ is very close to $u^{(n)}$, a perturbation of single point will send the optimal transport map evaluated at a particular $u_{i}\in u^{(n)}$ to infinity (see the construction of Figure 1). Thus, as the reference dataset becomes more disordered or random, the robustness of the optimal transport map increases.
$(0,0)$$(1,0)$$(-1,0)$$(0,1)$$(-1,0)$$(k,0)$$(k+1,0)$$(-1,0)$$(0,1)$$(-1,0)$
Figure 1: Example showing that the lower bound of Theorem 1.1 is sharp. The reference and target datasets $u^{(n)}$ and $X^{(n)}$ are both $\{(0,0),(1,0),(0,1),(-1,0),(0,-1)\}$. The Tukey depth of the origin is $3/5$ and the lower Tukey depth is $2/5$. Perturbing two points of $X^{(n)}$ as in the red dataset of the picture, we find that the optimal transport map from the blue the the red points maps $(0,0)$ to $(0,k)$, which escapes to the horizon. Therefore the breakdown point of the optimal transport map evaluated at the origin is $2/5$.
We now present our main theorem, which asserts that the breakdown point of $\partial\varphi=\partial\varphi_{\mu\to P}$, evaluated at $u\in\mathbb{R}^{d}$, corresponds to the Tukey depth of $u$ in the reference measure $\mu$. Recall that $\partial\varphi$ is the subdifferential of a l.s.c. convex function $\varphi$ whose gradient pushes $\mu$ forward to $P$.
Theorem 3.4.
Set $\mu,P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$. Let $\nabla\varphi$ be the optimal transport map pushing $\mu$ forward to $P$. Then the breakdown point of a canonical extension $\partial\varphi$ at a point $u$ with $\mathrm{TD}(u;\mu)>0$ is the Tukey depth of $u$ w.r.t. $\mu$, i.e.,
${\rm BP}(\partial\varphi(u),P)=\mathrm{TD}(u;\mu),\quad\text{irrespective of $\mu,P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$.}$
Remark 3.5.
After inspecting the proof of Theorem 3.4 (cf. 4.3 and 4.4), that a stronger notion of breakdown point leads to the same quantitative result. Namely, for every $u\in\mathbb{R}^{d}$ such that $\mathrm{TD}(u;\mu)>0$,
$\overline{\rm BP}(\partial\varphi(u),P)=\min\left\{\varepsilon\in(0,1):\sup_{\varphi_{\mu\to\nu}\in\Gamma_{\varepsilon}(P)}\sup_{v\in\partial\varphi_{\mu\to\nu}(u)}\|v\|=\infty\right\}=\mathrm{TD}(u;\mu).$
3.2 Breakdown points of the transport-based quantiles
We can now easily obtain the breakdown point of the empirical and population depth contours using the results from the previous subsection. Theorem 3.1 yields the breakdown points of the transport-based median and depth contours for discrete probability measures, while Theorem 3.4 gives the result for their continuous counterparts.
Corollary 3.6.
Let $X_{1},\dots,X_{n}$ be a sequence of vectors in $\mathbb{R}^{d}$ with all of the elements being different, and let $N=\max_{(v,u)\in\mathcal{S}^{d-1}\times\mathbb{R}^{d}}\left\{\sum_{i=1}^{n}\mathbf{1}_{\left(\langle v,u_{i}-u\rangle=0\right)}\right\}-1$. Then, irrespective of the dataset $X_{1},\dots,X_{n}$,
${\rm BP}(m_{\pm})\in[1/2-N/n,1/2],$
and for every $\tau\in[0,1/2)$ with ${\cal C}_{\pm}^{(n)}(\tau)\neq\emptyset$,
${\rm BP}({\cal C}_{\pm}^{(n)}(\tau))\in[\tau-N/n,\tau].$
Corollary 3.7.
Take a reference measure $\mu\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ such that $\max_{x\in\mathbb{R}^{d}}{\rm TD}(x;\mu))=1/2$. Then, for any $P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$,
${\rm BP}({m}_{\pm}({P}),P)=\frac{1}{2}$
and
${\rm BP}({\cal C}_{\pm}(\tau))=\tau\quad\text{holds for all $\tau\in(0,1/2)$}$
Remark 3.8.
From 3.6 and 3.7 we conclude that the reference measure does not have any influence on the population breakdown point as long as it attains the maximal Tukey depth of $1/2$. However, in the discrete case it might have an small impact as illustrated in 3.3.
4 Proof of main results
This section is devoted to the proofs of Theorem 3.1 and Theorem 3.4. Both arguments follow the same idea, which is based on separating a portion of the data into a ball escaping to the horizon in the direction given by $v_{u}\in\operatorname*{arg\,min}_{v\in\mathcal{S}^{d-1}}\mu\left(\{z:\langle v,z-u\rangle\geq 0\}\right)$. We will argue that a claimed breakdown point $\varepsilon^{*}$ holds by contradiction. We will show that if $\varepsilon^{*}$ is not the breakdown, then transport-based quantile function (or its empirical counterpart) would have to transform a convex cone into a non-convex cone that violates the mass-preserving property of transport maps. More precisely, we show that any monotone set-valued mapping $T:\mathbb{R}^{d}\to 2^{\mathbb{R}^{d}}$ transforms the convex cone
$\mathcal{C}_{x,e,\theta}=\{x+v\in\mathbb{R}^{d}:\ \langle v,e\rangle\geq\cos(\theta)\|v\|\},$ (5)
with $\|e\|=1$, into the non-convex cone
$\mathcal{A}_{u,e,\theta}=\left\{y\in\mathbb{R}^{d}:\langle u-y,e\rangle\leq\|u-y\|\frac{|\sin(\theta)|}{(1-\sin(\theta)^{2})^{1/2}}\right\}\quad\text{for}\ u=y^{x}\in T(x).$
We deliver a contradiction by showing that if $T$ pushes forward $\mu$ to $P$ (or its empirical measure) and $\varepsilon^{*}$ is not the breakdown, then the $P$-mass of $\mathcal{C}_{x,e,\theta}$ is larger than the $\mu$-mass of $\mathcal{A}_{u,e,\theta}$. Since $T(\mathcal{C}_{x,e,\theta})\subset\mathcal{A}_{u,e,\theta}$, this contradicts the mass-preserving property of the push forward $T$.
[…]
Figure 2: Diagram illustrating the strategy for the proof of Theorem 3.4. For the upper bound, we consider the mixture $(1-\varepsilon)P+\varepsilon Q_{k}$, where $Q_{k}$ is uniformly distributed on the ball $kv_{u}+\mathbb{B}$ for $v_{u}\in\operatorname*{arg\,min}_{v\in\mathcal{S}^{d-1}}\mu\left(\{z:\langle v,z-u\rangle\geq 0\}\right)$. If $\varepsilon>{\rm TD}(u;\mu)$ and a sequence $\{x_{k}\}_{k\in\mathbb{N}}$ of with $x_{k}\in\partial\varphi_{\varepsilon,k}(u):=\partial\varphi_{\mu\to(1-\varepsilon)P+\varepsilon Q_{k}}(u)$ remains bounded within a ball of radius $R$, then for sufficiently large $k$, the map $(\partial\varphi_{\varepsilon,k})^{-1}$ transforms the cone with vertex $x_{k}$ and angle $\theta_{k}$ into a set located to the right of the rays $r_{1}$ and $r_{2}$. As the ball escapes to the horizon, the angle $\theta_{k}$ tends to 00, causing the image of the ball to be contained within the left half-space defined by the hyperplane $H$, which is tangent to $v_{u}$. Consequently, as $\varepsilon>{\rm TD}(u;\mu)$, the $((1-\varepsilon)P+\varepsilon Q_{k})$-measure of the cone exceeds ${\rm TD}(u;\mu)$, while its image under $(\partial\varphi_{\varepsilon,k})^{-1}$ converges to a set with $\mu$-probability of ${\rm TD}(u;\mu)$. This leads to a contradiction. The strategy for the lower bound follows a similar approach.
The proof idea of the contradiction is illustrated in Figure 2. 4.1 formalizes that a monotone set-valued map $T$ transforms $\mathcal{C}_{x,e,\theta}$ into $\mathcal{A}_{u,e,\theta}$. Interested readers may note that similar results are commonly employed to show that a transport map between probability measures $\mu_{1}$ and $\mu_{2}$, where $\mu_{2}$ is supported on a strongly convex set, cannot map interior points of $\operatorname{supp}(\mu_{1})$ to boundary points of $\operatorname{supp}(\mu_{2})$ (cf. [11, 15]).
Lemma 4.1.
Let $T:{\rm dom}(T)\subset\mathbb{R}^{d}\to 2^{\mathbb{R}^{d}}$ be a monotone operator, i.e.,
$\langle x_{1}-x_{2},y_{1}-y_{2}\rangle\geq 0,\quad\text{for all }x_{1},x_{2}\in{\rm dom}(T),\ y_{1}\in T(x_{1})\ \text{and}\ y_{2}\in T(x_{2}).$ (6)
Set $x\in{\rm dom}(T)$. Then, for any $\theta\in(0,\pi/2)$ and any $e\in\mathcal{S}^{d-1}$, it holds that
$T(\mathcal{C}_{x,e,\theta}):=\bigcup_{z\in\mathcal{C}_{x,e,\theta}}T(z)\subset\mathcal{A}_{y^{x},e,\theta},$
for all $y^{x}\in T(x)$.
Proof.
Firstly, we construct an orthonormal basis $\{e_{1}\}_{i=1}^{d}$ with $e_{1}=e$. For a vector $v$ we denote as $(v_{1},\dots,v_{d})$ its coordinates w.r.t. the basis $\{e_{1}\}_{i=1}^{d}$. Fix $z\in\mathcal{C}_{x,e_{1},\theta}$. By definition of the cone, $z_{1}-x_{1}\geq\cos(\theta)\|z-x\|$. Hence,
$(z_{1}-x_{1})^{2}$ $\geq\cos(\theta)^{2}\|z-x\|^{2}$
$=\cos(\theta)^{2}\sum_{i=1}^{d}(z_{i}-x_{i})^{2}=(1-\sin(\theta)^{2})\sum_{i=1}^{d}(z_{i}-x_{i})^{2},$
which yields
$\sin(\theta)^{2}(z_{1}-x_{1})^{2}\geq(1-\sin(\theta)^{2})\sum_{i=2}^{d}(z_{i}-x_{i})^{2}.$
Since $\cos(\theta)\geq 0$, then $z_{1}-x_{1}\geq 0$, from where we can derive the bound
$\|(x-z)-(x_{1}-z_{1})e_{1}\|\leq\frac{|\sin(\theta)|}{(1-\sin(\theta)^{2})^{1/2}}(z_{1}-x_{1}).$ (7)
Fix $y^{z}\in T(z)$ and $y^{x}\in T(x)$. Then, by (6), it holds that $0\leq\langle y^{x}-y^{z},x-z\rangle$, so that
$0\leq\langle y^{x}-y^{z},x-z\rangle=\langle y^{x}-y^{z},x-z-(x_{1}-z_{1})e_{1}\rangle+(x_{1}-z_{1})\langle y^{x}-y^{z},e_{1}\rangle.$
Rearranging the terms and using Cauchy–Schwarz inequality we get
$(z_{1}-x_{1})\langle y^{x}-y^{z},e_{1}\rangle$ $\leq\langle y^{x}-y^{z},x-z-(x_{1}-z_{1})e_{1}\rangle$
$\leq\|y^{x}-y^{z}\|\|x-z-(x_{1}-z_{1})e_{1}\|.$
By (7), it holds that
$(z_{1}-x_{1})\langle y^{x}-y^{z},e_{1}\rangle\leq\|y^{x}-y^{z}\|\frac{|\sin(\theta)|}{(1-\sin(\theta)^{2})^{1/2}}(z_{1}-x_{1}).$
Since $z_{1}-x_{1}>0$ we can erase the term $z_{1}-x_{1}$ in both sides and the result follows. ∎
4.1 Population transport map: proof of Theorem 3.4
We start with a straightforward observation, which follows directly from [4, Theorem 4.1].
Lemma 4.2.
Set $\mu\in\mathcal{P}(\mathbb{R}^{d})$. Let $\{\theta_{k}\}_{k\in\mathbb{N}}\subset(0,\pi/2)$ and $\{\theta_{k}\}_{k\in\mathbb{N}}\subset(0,\pi/2)$ be such that $\theta_{k}\to 0^{+}$ and $v_{k}\to v$, respectively, as $k\to\infty$. Then
$\limsup_{k\to\infty}\mu(\mathcal{A}_{u,v_{k},\theta_{k}})\leq\mu(\{y:\langle y-u,v\rangle\geq 0\}).$
We will prove that ${\rm BP}(\partial\varphi(u),P)=\mathrm{TD}(u;\mu)$ by showing matching upper and lower bounds separately. Recall that for a probability measure $\nu$, the Borel map $\nabla\varphi_{\mu\to\nu}$ is the unique gradient of a convex function pushing $\mu$ forward to $\nu$. For the case $\nu=P$, we just write $\nabla\varphi=\nabla\varphi_{\mu\to P}$. 2.2 ensures that a point with (strictly) positive Tukey depth has always non-empty subdifferential. The following lemma shows the upper bound.
Lemma 4.3.
Let $\mu,P\in\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ and $Q_{k}$ be the Lebesgue uniform distribution on the ball $kv_{u}+\mathbb{B}$, $k\in\mathbb{N}$, for some
$v_{u}\in\operatorname*{arg\,min}_{v\in\mathcal{S}^{d-1}}\mu(\{y:\langle y-u,v\rangle\geq 0\}).$
Then, for every $\varepsilon>{\rm TD}(u;\mu)>0$, the sequence $\{\partial\varphi_{\varepsilon,k}(u)\}_{k\in\mathbb{N}}$, with $\partial\varphi_{\varepsilon,k}$ being any maximal monotone extension of $\nabla\varphi_{\mu\to P+\varepsilon(Q_{k}-P)}$, escapes to the horizon as $k\to+\infty$, i.e.,
$\inf_{y\in\partial\varphi_{\varepsilon,k}(u)}\|y\|\to+\infty.$
Proof.
Note first that as ${\rm TD}(u;\mu)>0$, 2.2 implies that $\partial\varphi_{\varepsilon,k}(u)\neq\emptyset$ for all $k\in\mathbb{N}$ and any maximal monotone extension of $\nabla\varphi_{\mu\to P+\varepsilon(Q_{k}-P)}$. We will show that for any sequence $\{x_{k}\}_{k\in\mathbb{N}}$, with $x_{k}\in\partial\varphi_{\varepsilon,k}(u)$ for all $k\in\mathbb{N}$, it must hold that $\|x_{k}\|\to\infty.$ To prove the claim we argue by contradiction. To do so, we can assume, after taking subsequences, that $x_{k}$ belongs to some ball $R\,\mathbb{B}$ for all $k\in\mathbb{N}$ and some finite $R>0$. Assume that $k\geq 2(R+1)$ and let $\mathcal{D}_{k}$ be the set of all rays arising from $x_{k}$ and touching $kv_{u}+\mathbb{B}$. Note that for all $k$ large enough there exists $e_{k}\in\mathcal{S}^{d-1}$ and $\theta_{k}\in(0,\pi/2)$ such that $\mathcal{D}_{k}=\mathcal{C}_{x_{k},e_{k},\theta_{k}}$ as defined in (5).
On the one hand, it holds that
$(P+\varepsilon(Q_{k}-P))(\mathcal{C}_{x_{k},e_{k},\theta_{k}})\geq\varepsilon Q_{k}(\mathcal{C}_{x_{k},e_{k},\theta_{k}})=\varepsilon>{\rm TD}(u;\mu)+\alpha$ (8)
for some $\alpha>0$. On the other hand, 4.1 and (2) imply that
$(\partial\varphi_{\varepsilon,k})^{-1}(\mathcal{C}_{x_{k},e_{k},\theta_{k}})\subset\mathcal{A}_{u,e_{k},\theta_{k}},$
which yields
$\mu((\partial\varphi_{\varepsilon,k})^{-1}(\mathcal{C}_{x_{k},e_{k},\theta_{k}}))<\mu(\mathcal{A}_{u,e_{k},\theta_{k}}).$ (9)
Note that $\theta_{k}\to 0$ and $e_{k}\to v_{u}$, due to $\|x_{k}\|\leq R$ and the fact that $kv_{u}+\mathbb{B}$ escapes to the horizon. The contradiction will be found by showing that as $k\to\infty$, the optimal transport map $\nabla\varphi_{\varepsilon,k}^{*}=(\nabla\varphi_{\varepsilon,k})^{-1}$ violates the push forward condition. Since $\partial\varphi_{\varepsilon,k}^{*}=(\partial\varphi_{\varepsilon,k})^{-1}=\{\nabla\varphi_{\varepsilon,k}^{*}(x)\}$ for $(P+\varepsilon(Q_{k}-P))$-a.e. $x$ and the Borel map $\nabla\varphi_{\varepsilon,k}^{*}$ pushes $P+\varepsilon(Q_{k}-P)$ forward to $\mu$, it must hold that
$\mu(\partial\varphi_{\varepsilon,k}^{*}(\mathcal{C}_{x_{k},e_{k},\theta_{n}}))=(P+\varepsilon(Q_{k}-P))(\mathcal{C}_{x_{k},e_{k},\theta_{k}}).$
We use (8) and (9) to get $\mu(\mathcal{A}_{u,e_{k},\theta_{k}})>{\rm TD}(u;\mu)+\alpha.$ Since $\theta_{k}\to 0$ and $e_{k}\to v_{u}$, 4.2 leads to the contradiction
${\rm TD}(u;\mu)$ $=\min_{v\in\mathcal{S}^{d-1}}\mu(\{y:\langle y-u,v\rangle\geq 0\})$
$=\mu(\{y:\langle y-u,v_{u}\rangle\geq 0\})=\lim_{k\to\infty}\mu(\mathcal{A}_{u,e_{k},\theta_{k}})>{\rm TD}(u;\mu)+\alpha.$
Therefore, the result follows. ∎
We now turn to the lower bound and argue again by contradiction. Assume that $\alpha={\rm TD}(u;\mu)-\varepsilon>0$ and that there exists a sequence $\{Q_{k}\}_{k\in\mathbb{N}}\subset\mathcal{P}^{a.c.}(\mathbb{R}^{d})$ such that $\|x_{k}\|\to\infty$ with $x_{k}\in\partial\varphi_{\varepsilon,k}(u)$ and $\partial\varphi_{\varepsilon,k}$ being any maximal monotone extension of $\nabla\varphi_{\mu\to P+\varepsilon(Q_{k}-P)}$ for all $k\in\mathbb{N}$. As $P$ is tight, for every $\beta>0$ there exists $R=R(\beta)>0$ such that $P(R\,\mathbb{B})\geq 1-\beta$. Now we argue as in the proof of 4.3 but with the cone $\mathcal{C}_{x_{k},e_{k},\theta_{k}}$ containing $R\,\mathbb{B}$ and having vertex $x_{k}$. As $\|x_{k}\|\to\infty$, the angle $\theta_{k}$ of the cone $\mathcal{C}_{x_{k},e_{k},\theta_{k}}$ tends to 00 and the direction $e_{k}$ tends (after taking subsequences) to some $v_{0}\in\mathcal{S}^{d-1}$. Then we obtain
$((1-\varepsilon)P+\varepsilon Q_{k})(\mathcal{C}_{x_{k},e_{k},\theta_{k}})$ $\geq(1-\varepsilon)P(R\,\mathbb{B})$
$\geq(1-({\rm TD}(u;\mu)-\alpha))(1-\beta)>1-{\rm TD}(u;\mu)+\alpha^{\prime},$
for some $\beta=\beta({\rm TD}(u;\mu),\alpha)$ small enough and $\alpha^{\prime}=\alpha^{\prime}(\beta,\alpha)>0$. Repeating exactly the same argument of 4.3 we obtain
$1-{\rm TD}(u;\mu)+\alpha^{\prime}$ $<((1-\varepsilon)P+\varepsilon Q_{k})(\mathcal{C}_{x_{k},e_{k},\theta_{k}})$
$=\mu((\partial\varphi_{\varepsilon,k})^{-1}(\mathcal{C}_{x_{k},e_{k},\theta_{k}}))\leq\mu(\mathcal{A}_{u,e_{k},\theta_{k}}).$
Hence, 4.2 yields $1-{\rm TD}(u;\mu)+\alpha^{\prime}\leq\mu(\{y:\langle y-u,v_{0}\rangle\geq 0\})$. However, as $\mu\ll\mathcal{L}_{d}$, it holds that
$\mu(\{y:\langle y-u,v_{0}\rangle\geq 0\})$ $\leq\max_{v\in\mathcal{S}^{d-1}}\mu(\{y:\langle y-u,v\rangle\geq 0\})$
$=\max_{v\in\mathcal{S}^{d-1}}(1-\mu(\{y:\langle y-u,v\rangle\leq 0\})$
$=1-{\rm TD}(u;\mu),$
which leads to a contradiction. Hence, the following result follows and the proof of Theorem 3.4 finishes here.
Lemma 4.4.
Let $u\in\mathbb{R}^{d}$ be such that ${\rm TD}(u;\mu)>0$. Then for every $\varepsilon\in[0,{\rm TD}(u;\mu))$, there exists $R>0$, such that
$\{x\in\partial\varphi_{\mu\to\nu}(u):\ \varphi_{\mu\to\nu}\in\Gamma_{\varepsilon}(P)\}\subset R\,\mathbb{B}.$
4.2 Empirical optimal transport: proof of Theorem 3.1
For two $n\times n$ real matrices $A=(A_{ij})$ and $B=(B_{ij})$, where $i,j\in[n]$ and $[n]=\{1,\dots,n\}$, let $\langle A,B\rangle_{F}=\sum_{i,j\in[n]}A_{ij}B_{ij}$ denote their Frobenius inner product. An optimal matching between the dataset $X^{(n)}$ and the reference points $u^{(n)}$ is any $n\times n$ real matrix $\gamma=(\gamma_{ij})$ solving
$\inf_{\gamma\in\Pi_{n}}\langle c,\pi\rangle.$
where $c=(c_{ij})$ is a symmetric $n\times n$ cost matrix with elements $c_{ij}=(\|u_{i}-X_{j}\|^{2})$, and $\Pi_{n}$ denotes the Birkhoff polytope of doubly stochastic matrices [5] defined as
$\Pi_{n}=\left\{\pi\in R^{n\times n}:\,\sum_{i=1}^{n}\pi_{ij}=1,\;\sum_{j=1}^{n}\pi_{ij}=1,\;\pi_{ij}\geq 0\right\}.$
A celebrated theorem due to Birkhoff states that the vertexes of $\Pi_{n}$ is the set of permutation matrices. This implies that an optimal matching can be identified with an optimal permutation $\hat{\sigma}$ on the set $\{1,\dots,n\}$. Therefore, an empirical optimal transport map $\hat{T}_{n}=\hat{T}_{X^{(n)}}$ can be described as $\hat{T}_{n}(u_{i})=X_{\hat{\sigma}(i)}$, where $\hat{\sigma}$ is an optimal permutation, which in general might not be unique. However, it is well-known that
$\langle u_{i}-u_{j},\hat{T}_{n}(u_{i})-\hat{T}_{n}(u_{j})\rangle\geq 0\quad\text{ for all }i,j\in\{1,\dots,n\}$ (10)
for any optimal transport map $\hat{T}_{n}$. The monotonicity property (10) is the key ingredient necessary to apply 4.1 in the proof of Theorem 3.1. The claimed finite sample breakdown point results follow from the upper and lower bounds that we establish in the following two lemmas. Their proofs follow along the lines of those of given for 4.3 and 4.4.
Lemma 4.5.
Let $\ell\in\{1,\dots,n\}$ and $X^{(n,\ell)}_{k}=\{X_{1}\dots,X_{n-\ell},v_{1}^{(k)},\dots,v_{\ell}^{(k)}\}$ where $\{v_{1}^{(k)},\dots,v_{\ell}^{(k)}\}\subset kv_{u}+\mathbb{B}$, $k\in\mathbb{N}$ for some
$v_{u}\in\operatorname*{arg\,min}_{v\in\mathcal{S}^{d-1}}\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\left(\langle v,u_{i}-u\rangle\geq 0\right)}\right\}.$
Then, for any $u_{i}\in\mathbb{R}^{d}$ be such that ${\rm TD}(u_{i};u_{1},\dots,u_{n})<\ell/n$, it holds that
$\lim_{k\to\infty}\|\hat{T}_{X^{(n,\ell)}_{k}}(u_{i})\|=\infty.$
Proof.
We argue by contradiction. After taking a subsequence, we can assume that $\{\hat{T}_{X^{(n,\ell)}_{k}}(u_{i})\}_{k\in\mathbb{N}}$ is contained in $R\,\mathbb{B}$ for some $R>0$. Let $\mathcal{C}_{x_{k},e_{k},\theta_{k}}$ be the cone of rays arising from $\hat{T}_{X^{(n,\ell)}_{k}}(u_{i})$ and touching the set $\{v_{1}^{(k)},\dots,v_{\ell}^{(k)}\}$. It holds that
$\nu_{\ell,k}(\mathcal{C}_{x_{k},e_{k},\theta_{k}})\geq\frac{\ell}{k}>{\rm TD}(u_{i};u^{(n)})+\alpha$ (11)
for some $\alpha>0$, where $\nu_{\ell,k}=\frac{1}{n}\sum_{x\in X^{(n,\ell)}_{k}}\delta_{x}$ denotes empirical measure of the dataset $X^{(n,\ell)}_{k}$. Since $\hat{T}_{X^{(n,\ell)}_{k}}$ is monotone (cf.  (10)), 4.1 implies that
$(\hat{T}_{X^{(n,\ell)}_{k}})^{-1}(X^{(n,\ell)}_{k}\cap\mathcal{C}_{x_{k},e_{k},\theta_{k}})\subset\mathcal{A}_{u_{i},e_{k},\theta_{k}},$
from where we deduce that
$\begin{split}{\rm TD}(u_{i};u^{(n)})+\alpha&<\nu_{\ell,k}(\mathcal{C}_{x_{k},e_{k},\theta_{k}})\\ &=\mu_{n}((\hat{T}_{X^{(n,\ell)}_{k}})^{-1}(X^{(n,\ell)}_{k}\cap\mathcal{C}_{x_{k},e_{k},\theta_{k}})<\mu_{n}(\mathcal{A}_{u_{i},e_{k},\theta_{k}}),\end{split}$ (12)
where $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{u_{i}}$. Hence, 4.2 leads to a contradiction when taking limits in (12) as $k\to\infty$. (Note that, as in 4.3, $\theta_{k}\to 0$ and $e_{k}\to v$.) ∎
Lemma 4.6.
Fix $i\in\{1,\dots,n\}$. Then, for every $\ell\in\mathbb{N}$ such that ${\rm TD}^{-}(u_{i};u^{(n)})>\ell/n$, there exists a radius $R>0$ such that $\hat{T}_{X^{(n,\ell)}_{k}}(u_{i})\in R\,\mathbb{B}$ for all $X^{(n,\ell)}_{k}\in\mathcal{Q}_{\ell,n}$.
Proof.
Let $R>0$ be such that $X_{i}\in R\,\mathbb{B}$ for all $i=1,\dots,n$. Assume by contradiction that there exists a sequence $\{X^{(n,\ell)}_{k}\}_{k\in\mathbb{N}}\subset\mathcal{Q}_{\ell,n}$ such that $\|\hat{T}_{X^{(n,\ell)}_{k}}(u_{i})\|\to+\infty$ as $k\to+\infty$. We define the counting measures $\nu_{\ell,k}=\frac{1}{n}\sum_{u\in X^{(n,\ell)}_{k}}\delta_{u}$ and $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{u}$. We repeat the same steps of the proof of 4.4 but exchanging $P+\varepsilon(Q-P)$ by $\nu_{\ell,k}$ and $\mu$ by $\mu_{n}$ to derive the existence of $\alpha>0$ such that
$\frac{n+1}{n}-{\rm TD}^{-}(u_{i};u^{(n)})+\alpha$ $\leq\mu_{n}(\{y:\langle y-u,v\rangle\geq 0\})$
$\leq\max_{v\in\mathcal{S}^{d-1}}\mu_{n}(\{y:\langle y-u,v\rangle\geq 0\}),$
which leads to a contradiction. ∎